# Estimating Dense-Packed Zone Height in Liquid-Liquid Separation: A Physics-Informed Neural Network Approach

### v1

@@ 4 Physics-informed neural network model of the settler dynamics
In this section, we present the dynamic modeling of the gravity settler through the PINN approach. In Section˜4.1, we present the PINN architecture. Section˜4.2 discusses the PINN training strategy, Section˜4.3 presents the purely data-driven neural network for comparison and Section˜4.4 presents prediction results and comparison of different models.
@@ 4.2 Training strategy
In the second stage, we fine-tune the pretrained models using data from the experimental setup, $\mathcal{D}_{\mathrm{exp}}$. To prevent the model from forgetting what it has learned during pretraining, we reduce the learning rate and the number of training epochs. The overall training strategy is depicted in Fig.˜8.
Note that the immeasurable internal flows $\mathbf{z}(t)=[Q_{\mathrm{c}}(t),Q_{\mathrm{s}}(t)]^{\top}$ are outputs of the mechanistic model. We therefore use these quantities as training data during the pretraining stage. During fine-tuning, however, their corresponding loss terms are disabled, as these internal flows are not directly measurable and no experimental data are available for them.
The associated semi-empirical submodels, which define $Q_{\mathrm{c}}(t)$ and $Q_{\mathrm{s}}(t)$ as functions of the system states (cf. Eq.˜2), are not embedded directly into the PINN. These submodels are lumped formulations obtained by first resolving discretized internal compartments and subsequently aggregating their contributions. In particular, the separator is discretized along the axial direction into 200 segments and across 50 droplet diameter classes, resulting in approximately 200 × 50 = 10,000 internal compartments.
Embedding these calculations into the PINN would either require introducing a large number of additional outputs to represent the states of the discretized compartments or evaluating the submodels sequentially using PINN-predicted states during training. In both cases, this would lead to a prohibitive computational cost, as the full set of discretized compartments would need to be evaluated repeatedly at each training iteration. Instead, we employ the semi-empirical submodels offline to generate a fixed dataset of trajectories for pretraining. This dataset is generated once and reused throughout training, yielding a computationally efficient procedure while still incorporating the information provided by the sedimentation and coalescence submodels.
@@
$MSE_{total}$ $=\lambda_{1}MSE_{data}+\lambda_{2}MSE_{physics}+MSE_{init},$ (4a)
[…]
@@
$\begin{split}MSE_{physics}&=\frac{1}{3|\mathcal{D}_{\mathrm{physics}}|}\sum_{j=1}^{|\mathcal{D}_{\mathrm{physics}}|}\Biggl[\lambda_{g}\left(\bm{g}_{\mathrm{sep}}\left(Q_{\textrm{in}}(t_{j}),\hat{Q}_{\textrm{bot}}(t_{j}),\hat{Q}_{\textrm{top}}(t_{j})\right)\right)^{2}\\ &+\left(\dot{\hat{h}}_{\textrm{HP}}(t_{j})-\bm{f}_{\mathrm{DP}}\left(\hat{h}_{\textrm{HP}}(t_{j}),\hat{h}_{\textrm{DP}}(t_{j}),Q_{\textrm{in}}(t_{j}),\hat{Q}_{\textrm{bot}}(t_{j}),\hat{Q}_{c}(t_{j}),\hat{Q}_{s}(t_{j})\right)\right)^{2}\\ &+\left(\dot{\hat{h}}_{\textrm{DP}}(t_{j})-\bm{f}_{\mathrm{HP}}\left(\hat{h}_{\textrm{HP}}(t_{j}),\hat{h}_{\textrm{DP}}(t_{j}),Q_{\textrm{in}}(t_{j}),\hat{Q}_{\textrm{bot}}(t_{j}),\hat{Q}_{c}(t_{j}),\hat{Q}_{s}(t_{j})\right)\right)^{2}\Biggr]\end{split}$ (4c)
@@
$MSE_{init}$ $=\frac{1}{2|\mathcal{D}_{\mathrm{init}}|}\sum_{j=1}^{|\mathcal{D}_{\mathrm{init}}|}\Biggl[\left(\hat{h}_{\textrm{DP}_{j}}(t=0)-h_{\textrm{DP}_{j}}(0)\right)^{2}-\left(\hat{h}_{\textrm{HP}_{j}}(t=0)-h_{\textrm{HP}_{j}}(0)\right)^{2}\Biggr]$ (4d)
@@
Here, $MSE_{\mathrm{data}}$ corresponds to the loss term accounting for the measurement data, $MSE_{\mathrm{physics}}$ corresponds to the loss term computed based on the available physics knowledge, and $MSE_{\mathrm{init}}$ measures the mismatch between the NN predictions at $t=0$ and the initial values $h_{\textrm{HP}_{j}}(0)$ and $h_{\textrm{DP}_{j}}(0)$. The subscript $j$ indexes a finite set of samples taken at times $t_{j}$, each associated with initial values $h_{\textrm{HP}_{j}}(0)$ and $h_{\textrm{DP}_{j}}(0)$ and control action $Q_{in,j}$; for simplicity, initial values and control action are omitted from the notation. $\lambda_{1}$ and $\lambda_{2}$ represent the weights assigned to the data and physics condition loss terms, respectively, $\lambda_{g}$ establishes a weighting between the algebraic and the differential equations within the physics loss term, and $\lambda_{z}$ establishes a weighting between the measurable and immeasurable PINN outputs within the data loss term. $\mathcal{D}_{\mathrm{data}}$ refers to the dataset consisting of measurements; specifically, we use $\mathcal{D}_{\mathrm{sim}}$ in the pretraining stage and $\mathcal{D}_{\mathrm{exp}}$ in the fine-tuning stage. The simulation dataset $\mathcal{D}_{\mathrm{sim}}$ contains the internal flows $Q_{\mathrm{c}}$ and $Q_{\mathrm{s}}$, yielding $n_{\mathrm{out}}=6$, where $n_{\mathrm{out}}$ denotes the number of PINN outputs. These quantities are absent from the experimental dataset $\mathcal{D}_{\mathrm{exp}}$; consequently, fine-tuning is performed with $n_{\mathrm{out}}=4$ and the corresponding loss weight for immeasurable flows $\lambda_{z}$ is set to zero.
@@
We employ a two-optimizer training scheme for the PINN in each training stage. We start the training with a stochastic gradient descent (SGD) based optimizer (ADAM) (kingma2017adammethodstochasticoptimization), then switch to Limited-memory Broyden–Fletcher–Goldfarb–Shanno algorithm (L-BFGS) (liu1989limited) for refinement. L-BFGS has been shown to be a very successful optimizer for training PINNs (Markidis2021TheSolvers), however, since it is a local optimizer, it has a tendency to get stuck in local minima (Markidis2021TheSolvers). du2019gradientdescentfindsglobal have shown that SGD, due to its stochastic directions taken in optimization landscape can lead to (near) global training. Thus, we first use SGD to mitigate getting stuck in a local minimum, then we use L-BFGS to refine the network weights. In the pretraining stage, we use a fixed number of training epochs for each optimizer, 2000 for ADAM with a fixed learning rate of 0.001 and 300 for L-BFGS with the strong-Wolfe condition check activated. Moreover, we apply inverse Dirichlet weighting (IDW) (Maddu2022InverseNetworks) every five epochs during the ADAM optimization stage to dynamically update the individual loss term weights, $\lambda_{1}$, $\lambda_{2}$, $\lambda_{g}$ and $\lambda_{z}$ in Eq. (4). In the fine-tuning stage, we reduce the learning rate by an order of magnitude to 0.0001 and lower the number of training epochs to 1000 for the ADAM optimizer and 200 for the L-BFGS optimizer. Moreover, we disable the IDW as the weightings of the different loss components are kept identical to the final values established in pretraining.
@@ 4.3 Benchmark vanilla neural network model
As a comparison for the PINN, we implement a benchmark vanilla neural network (VNN) model. The VNN has the same input configuration as the PINN: it takes time $t$, the current heights $\mathbf{x}(t=0)=[h_{\textrm{DP}}(0),h_{\textrm{HP}}(0)]^{\top}$ and control inputs $\mathbf{u}=[Q_{\textrm{in}}]$ as inputs and predicts the corresponding measurable heights $h_{\textrm{DP}}(t)$, $h_{\textrm{HP}}(t)$ and outlet flows $Q_{\textrm{bot}}(t)$, $Q_{\textrm{top}}(t)$. Note that the VNN cannot predict the immeasurable internal flows $Q_{\mathrm{c}}(t)$ and $Q_{\mathrm{s}}(t)$, since no experimental data can be gathered for those. The VNN is trained in a data-driven fashion, i.e., without the physics-based loss term in Eq.˜4c. For comparison with the PINN model, we adopt the same network architecture as used for the PINN in Section˜4.2, i.e, the same number of layers and neurons and the same activation functions. Moreover, we use the same datasets, except for, $\mathcal{D}_{\mathrm{physics}}$, which is used exclusively in physics-based training.
@@ 4.4 Prediction results
We analyze how the models predict the experimental trajectories of phase heights when initialized from a known state at $\tau=0$. We compare the accuracies of the PINN, VNN, and mechanistic model predictions, and examine the effects of pretraining. For both PINN and VNN, we employ an ensemble of models. Using an ensemble enables averaging over predictions to mitigate the impact of outlier models and provides a means to quantify prediction uncertainty (dietterich2000ensemble). The ensemble size for both PINN and VNN is fixed at $n_{\mathrm{ensemble}}=40$.
@@
In Figs. 9(a) and 9(b), we present the results for the interpolation trajectory for the heavy-phase height and the DPZ height, respectively. While the two-stage trained PINN ensemble tracks the averaged heights with good accuracy, the VNN ensemble mean overshoots with regard to the DPZ height. For both model types, the non-pretrained versions yield noticeably poorer predictions, with the effect being especially pronounced for the VNN. Moreover, we see that the mechanistic model predictions diverge, predicting flooding of the separator with the dispersed-phase at around $\tau=$380\text{\,}\mathrm{s}\text{/}$\$.
In Figs. 10(a) and 10(b), we present the extrapolation trajectory results. The overall trends are similar to those observed for the interpolation trajectories: The two-stage trained PINN ensemble provides a fairly accurate tracking (albeit a slight overshoot towards the end), the VNN and the non-pretrained models perform significantly worse and the mechanistic model predicts flooding at around $\tau=$420\text{\,}\mathrm{s}\text{/}$\$.
@@
Figure 9: Simulation results for the interpolation test trajectory. Predictions from an individual model are shown in transparent red, whereas the ensemble mean estimation is shown with a red dashed line.
@@
Figure 10: Simulation results for the extrapolation test trajectory. Predictions from an individual model are shown in transparent red, whereas the ensemble mean estimation is shown with a red dashed line.

### v2

@@ 4 Physics-informed neural network model of the settler dynamics
In this section, we present the dynamic modeling of the gravity settler through the PINN approach. In Section˜4.1, we present the PINN architecture. The physics equations and the associated modeling assumptions are presented in Section˜4.2 and data handling for the PINN training is presented in Section˜4.3. Section˜4.4 discusses the PINN training strategy, Section˜4.5 presents the purely data-driven neural network for comparison, and Section˜4.6 presents prediction results and comparison of different models.
@@ 4.1 Model architecture
4.2 Physics equations and modeling assumptions
The physics-informed component of the proposed PINN is based on the volume-balance equations governing the phase height dynamics in the gravity settler, as introduced in Section˜3. In the present work, only these macroscopic balance equations are incorporated into the PINN as soft constraints, while the detailed droplet coalescence and sedimentation submodels are not embedded.
Based on the volume balance equations in Eq.˜1, the physics constraints are formulated as residuals that penalize deviations from the underlying dynamics. Specifically, we define an algebraic residual ($e_{1}$) describing the overall flow balance of the separator, and differential residuals ($e_{2}$ and $e_{3}$) describing the phase height dynamics:
$e_{1}(t)$ $=g_{\mathrm{sep}}\big(Q_{\mathrm{in}}(t),\hat{Q}_{\mathrm{bot}}(t),\hat{Q}_{\mathrm{top}}(t)\big),$ (4a)
$e_{2}(t)$ $=\dot{\hat{h}}_{\mathrm{DP}}(t)-f_{\mathrm{DP}}\big(\hat{h}_{\mathrm{DP}}(t),\hat{h}_{\mathrm{HP}}(t),Q_{\mathrm{in}}(t),\hat{Q}_{\mathrm{bot}}(t),\hat{Q}_{\mathrm{c}}(t),\hat{Q}_{\mathrm{s}}(t)\big),$ (4b)
$e_{3}(t)$ $=\dot{\hat{h}}_{\mathrm{HP}}(t)-f_{\mathrm{HP}}\big(\hat{h}_{\mathrm{HP}}(t),Q_{\mathrm{in}}(t),\hat{Q}_{\mathrm{bot}}(t),\hat{Q}_{\mathrm{c}}(t),\hat{Q}_{\mathrm{s}}(t)\big)$ (4c)
Here, $f_{\mathrm{DP}}$ and $f_{\mathrm{HP}}$ denote the right-hand sides of the volume-balance equations defined in Section˜3. The time derivatives of the predicted phase heights, $\dot{\hat{h}}$, are obtained via automatic differentiation of the neural network outputs with respect to the input time. These residuals define the physics-informed component of the training objective and penalize deviations from the governing equations.
Restricting the PINN to the volume-balance equations is mainly motivated by computational considerations. The semi-empirical submodels that define $Q_{\mathrm{c}}(t)$ and $Q_{\mathrm{s}}(t)$ as functions of the system states (cf. Eq.˜2) are lumped formulations obtained by resolving discretized internal compartments and subsequently aggregating their contributions. In particular, the separator is discretized along the axial direction into 200 segments and across 50 droplet diameter classes, resulting in approximately 200 × 50 = 10,000 internal compartments. Embedding these calculations into the PINN would either require introducing a tremendous number of additional outputs to represent the states of the discretized compartments or evaluating the submodels sequentially using PINN-predicted states during training. In both cases, this would lead to a prohibitive computational cost, as the full set of discretized compartments would need to be evaluated repeatedly at each training iteration. Moreover, the internal states associated with these submodels are not measurable in the experimental setup.
Instead of embedding the submodels for $Q_{\mathrm{c}}(t)$ and $Q_{\mathrm{s}}(t)$ into the PINN, the influence of these submodels is incorporated indirectly through the synthetic data used during pretraining, which is generated using the full mechanistic model. In this way, the PINN may learn an implicit representation of the underlying physics while maintaining a tractable model structure.
This modeling choice introduces a trade-off between physical completeness and computational efficiency. While the reduced physics representation may limit extrapolation in regimes dominated by droplet-scale phenomena, the combination of physics-informed training and experimental fine-tuning enables accurate phase height estimation in the operating conditions considered, as will be shown in Section˜5.3.
4.3 Data handling for PINN training
Based on the preprocessed experimental dataset $\mathcal{D}_{\mathrm{exp}}$ described in Section˜2.4, phase height and flow-rate measurements are available at a uniform sampling interval of $1\text{\,}\mathrm{s}\text{/}$. This results in 789 segments (i.e., training samples), each corresponding to a single-step transition over $1\text{\,}\mathrm{s}\text{/}$. To align with the PINN formulation, which is defined over a temporal domain of $t\in[0,1]\,$\mathrm{s}\text{/}$\$, the trajectories are partitioned into consecutive segments of duration $1\text{\,}\mathrm{s}\text{/}$. Each segment represents a single-step state transition from process time $\tau_{k}$ to $\tau_{k+1}=\tau_{k}+$1\text{\,}\mathrm{s}\text{/}$\$ and contains two measurement points for the phase heights, corresponding to the segment boundaries at $t=0$ and $t=1$.
For each segment, the PINN takes as input the initial states, the applied control input, and a specified value of the continuous time variable $t\in[0,1]$, and returns the corresponding system states and outputs at that time. The explicit inclusion of time as an input enables the use of automatic differentiation to evaluate time derivatives of the network outputs. At the same time, this formulation provides a continuous-time representation of the system dynamics within each segment, allowing the PINN to generate predictions at arbitrary intermediate times.
An analogous procedure is applied to the simulation dataset $\mathcal{D}_{\mathrm{sim}}$ generated from the mechanistic model described in Section˜3. Specifically, we generate 1000 segments of duration $1\text{\,}\mathrm{s}\text{/}$, with inputs sampled within the bounds of the experimental data (cf. Tab.˜3). Within each segment, the mechanistic model is evaluated at a finer temporal resolution of $\Delta t=$0.1\text{\,}\mathrm{s}\text{/}$\$, resulting in additional training points that enhance the representation of the system dynamics over the interval.
It is important to note that, while the PINN yields continuous-time predictions, the experimental data provide measurements only at the discrete sampling instants corresponding to the segment boundaries. Intermediate predictions within each segment are therefore not directly supervised by experimental data, but are supported during training by both the pretraining dataset $\mathcal{D}_{\mathrm{sim}}$, which provides more densely sampled data, and the collocation points in $\mathcal{D}_{\mathrm{physics}}$. The validation and test trajectories are not included in $\mathcal{D}_{\mathrm{exp}}$ and are used exclusively for model evaluation.
@@ 4.4 Training strategy
In the second stage, we fine-tune the pretrained models using data from the experimental setup, $\mathcal{D}_{\mathrm{exp}}$. The experimental dataset used for fine-tuning includes measurements of both phase heights and flow rates. To prevent the model from forgetting what it has learned during pretraining, we reduce the learning rate and the number of training epochs. The overall training strategy is depicted in Fig.˜8.
Note that the immeasurable internal flows $\mathbf{z}(t)=[Q_{\mathrm{c}}(t),Q_{\mathrm{s}}(t)]^{\top}$ are outputs of the mechanistic model. We therefore use these quantities as training data during the pretraining stage. During fine-tuning, however, their corresponding loss terms are disabled, as these internal flows are not measurable and thus no experimental data are available for them.
@@
$MSE_{total}$ $=\lambda_{\mathrm{data}}MSE_{\mathrm{data}}+\lambda_{\mathrm{latent}}MSE_{\mathrm{latent}}+\lambda_{\mathrm{init}}MSE_{\mathrm{init}}$ (5a)
$+\lambda_{\mathrm{diff}}MSE_{\mathrm{diff}}+\lambda_{\mathrm{alg}}MSE_{\mathrm{alg}},$ (5b)
@@
$\begin{split}MSE_{\mathrm{data}}&=\frac{1}{4|\mathcal{D}_{\mathrm{data}}|}\sum_{j=1}^{|\mathcal{D}_{\mathrm{data}}|}\Biggl[\left(\hat{h}_{\textrm{DP}}(t_{j})-h_{\textrm{DP}}(t_{j})\right)^{2}+\left(\hat{h}_{\textrm{HP}}(t_{j})-h_{\textrm{HP}}(t_{j})\right)^{2}\\ &+\left(\hat{Q}_{\textrm{bot}}(t_{j})-Q_{\textrm{bot}}(t_{j})\right)^{2}+\left(\hat{Q}_{\textrm{top}}(t_{j})-Q_{\textrm{top}}(t_{j})\right)^{2}\Biggr]\end{split}$ (5c)
@@
$MSE_{\mathrm{latent}}$ $=\frac{1}{2|\mathcal{D}_{\mathrm{data}}|}\sum_{j=1}^{|\mathcal{D}_{\mathrm{latent}}|}\Biggl[\left(\hat{Q}_{\mathrm{c}}(t_{j})-Q_{\mathrm{c}}(t_{j})\right)^{2}+\left(\hat{Q}_{\mathrm{s}}(t_{j})-Q_{\mathrm{s}}(t_{j})\right)^{2}\Biggr]$ (5d)
$MSE_{\mathrm{init}}$ $=\frac{1}{2|\mathcal{D}_{\mathrm{init}}|}\sum_{j=1}^{|\mathcal{D}_{\mathrm{init}}|}\Biggl[\left(\hat{h}_{\textrm{DP}_{j}}(t=0)-h_{\textrm{DP}_{j}}(0)\right)^{2}-\left(\hat{h}_{\textrm{HP}_{j}}(t=0)-h_{\textrm{HP}_{j}}(0)\right)^{2}\Biggr]$ (5e)
$MSE_{\mathrm{diff}}$ $=\frac{1}{2|\mathcal{D}_{\mathrm{physics}}|}\sum_{j=1}^{|\mathcal{D}_{\mathrm{physics}}|}\Biggl[\left(e_{2}(t_{j})\right)^{2}+\left(e_{3}(t_{j})\right)^{2}\Biggr]$ (5f)
$MSE_{\mathrm{algebraic}}$ $=\frac{1}{|\mathcal{D}_{\mathrm{physics}}|}\sum_{j=1}^{|\mathcal{D}_{\mathrm{physics}}|}(e_{1}(t_{j}))^{2}$ (5g)
@@
Here, $MSE_{\mathrm{data}}$ denotes the loss term associated with measurement data for phase heights and outlet flow rates, while $MSE_{\mathrm{latent}}$ corresponds to the loss for the latent internal flow rates. $MSE_{\mathrm{init}}$ quantifies the mismatch between the neural network predictions at $t=0$ and the prescribed initial values $h_{\mathrm{HP}_{j}}(0)$ and $h_{\mathrm{DP}_{j}}(0)$. The terms $MSE_{\mathrm{diff}}$ and $MSE_{\mathrm{alg}}$ represent the physics-based loss contributions derived from the differential and algebraic residuals defined in Eq.˜4. The subscript $j$ indexes a finite set of samples taken at times $t_{j}$, each associated with initial values $h_{\textrm{HP}_{j}}(0)$ and $h_{\textrm{DP}_{j}}(0)$ and control action $Q_{in,j}$; for simplicity, initial values and control action are omitted from the notation. $\lambda_{\mathrm{data}}$, $\lambda_{\mathrm{latent}}$, $\lambda_{\mathrm{diff}}$, $\lambda_{\mathrm{alg}}$, and $\lambda_{\mathrm{init}}$ denote the weights assigned to the data, latent, differential physics, algebraic physics, and initial condition loss terms, respectively. $\mathcal{D}_{\mathrm{data}}$ refers to the dataset consisting of measurements; specifically, we use $\mathcal{D}_{\mathrm{sim}}$ in the pretraining stage and $\mathcal{D}_{\mathrm{exp}}$ in the fine-tuning stage. The simulation dataset $\mathcal{D}_{\mathrm{sim}}$ contains the internal flows $Q_{\mathrm{c}}(t)$ and $Q_{\mathrm{s}}(t)$. These quantities are absent from the experimental dataset $\mathcal{D}_{\mathrm{exp}}$; consequently, fine-tuning is performed with the corresponding loss weight for the latent internal flows set to zero ($\lambda_{\mathrm{latent}}$ = 0) . Note that the samples used to compute $MSE_{\mathrm{data}}$ and $MSE_{\mathrm{latent}}$ are drawn from the same dataset $\mathcal{D}_{\mathrm{data}}$, and similarly the samples used for $MSE_{\mathrm{diff}}$ and $MSE_{\mathrm{alg}}$ are both drawn from $\mathcal{D}_{\mathrm{physics}}$.
@@
We employ a two-optimizer training scheme for the PINN in each training stage. We start the training with a stochastic gradient descent (SGD) based optimizer (ADAM) (kingma2017adammethodstochasticoptimization), then switch to Limited-memory Broyden–Fletcher–Goldfarb–Shanno algorithm (L-BFGS) (liu1989limited) for refinement. L-BFGS has been shown to be a very successful optimizer for training PINNs (Markidis2021TheSolvers), however, since it is a local optimizer, it has a tendency to get stuck in local minima (Markidis2021TheSolvers). du2019gradientdescentfindsglobal have shown that SGD, due to its stochastic directions taken in optimization landscape can lead to (near) global training. Thus, we first use SGD to mitigate getting stuck in a local minimum, then we use L-BFGS to refine the network weights. In the pretraining stage, we use a fixed number of training epochs for each optimizer, 2000 for ADAM with a fixed learning rate of 0.001 and 300 for L-BFGS with the strong-Wolfe condition check activated. Moreover, we apply inverse Dirichlet weighting (IDW) (Maddu2022InverseNetworks) every five epochs during the ADAM optimization stage to dynamically update the individual loss term weights, $\lambda_{\mathrm{data}}$, $\lambda_{\mathrm{diff}}$, $\lambda_{\mathrm{init}}$, $\lambda_{\mathrm{alg}}$ and $\lambda_{\mathrm{latent}}$ in Eq. (5). We note that IDW is adopted here as a practical training strategy based on preliminary empirical observations of improved stability and performance compared to a static weighting, consistent with its use in our previous work (VELIOGLU2025108899); a systematic quantitative evaluation of alternative weighting schemes is beyond the scope of this study. We report the evolution of the adaptive loss weights (IDW) and the corresponding weighted loss terms during pre-training in Fig.˜9. Although the weights span several orders of magnitude, the weighted loss contributions are of comparable magnitude. In the fine-tuning stage, we reduce the learning rate by an order of magnitude to 0.0001 and lower the number of training epochs to 1000 for the ADAM optimizer and 200 for the L-BFGS optimizer. Moreover, we disable the IDW as the weightings of the different loss components are kept identical to the final values established in pretraining.
(a) IDW weights
(b) Weighted loss terms
Figure 9: Evolution of adaptive loss weights (IDW) and corresponding weighted loss terms during pre-training, averaged over the ensemble. During fine-tuning, the weights are fixed to the final values from the pre-training.
@@ 4.5 Benchmark vanilla neural network model
As a comparison for the PINN, we implement a benchmark vanilla neural network (VNN) model. The VNN has the same input configuration as the PINN: it takes time $t$, the current heights $\mathbf{x}(t=0)=[h_{\textrm{DP}}(0),h_{\textrm{HP}}(0)]^{\top}$ and control inputs $\mathbf{u}=[Q_{\textrm{in}}]$ as inputs and predicts the corresponding measurable heights $h_{\textrm{DP}}(t)$, $h_{\textrm{HP}}(t)$ and outlet flows $Q_{\textrm{bot}}(t)$, $Q_{\textrm{top}}(t)$. Note that the VNN cannot predict the immeasurable internal flows $Q_{\mathrm{c}}(t)$ and $Q_{\mathrm{s}}(t)$, since no experimental data can be gathered for those. The VNN is trained in a data-driven fashion, i.e., without the physics-based loss terms in Eq.˜5f and Eq.˜5g. For comparison with the PINN model, we adopt the same network architecture as used for the PINN in Section˜4.4, i.e, the same number of layers and neurons and the same activation functions. Moreover, we use the same datasets, except for, $\mathcal{D}_{\mathrm{physics}}$, which is used exclusively in physics-based training.
@@ 4.6 Prediction results
We analyze how the models predict the experimental trajectories of phase heights when initialized from a known state at $\tau=0$. We compare the accuracies of the PINN, VNN, and mechanistic model predictions, and examine the effects of pretraining. For both PINN and VNN, we employ an ensemble of models. Using an ensemble enables averaging over predictions to mitigate the impact of outlier models and provides a means to quantify prediction uncertainty (dietterich2000ensemble). The ensemble size for both PINN and VNN is fixed at $n_{\mathrm{ensemble}}=40$. To construct the ensemble, we train $n_{\mathrm{ensemble}}=40$ models independently with identical architectures and training data, but different random initialization of the network parameters. Due to the nonconvex loss landscape and stochasticity of the training algorithm, this may result in different trained parameter configurations and, consequently, a distribution of predictions. The ensemble spread is interpreted as an approximation of epistemic uncertainty arising from limited data and training variability.
We report prediction results in Fig.˜10 and Fig.˜11. To quantitatively assess model performance, we report the root mean square error (RMSE) of the ensemble-mean predictions over each trajectory, defined as
$\mathrm{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\hat{y}_{i}-y_{i}\right)^{2}},$
where $\hat{y}_{i}$ and $y_{i}$ denote the predicted and reference values, respectively, and $N$ is the number of samples. In addition, to quantify prediction uncertainty, we report the time-averaged interquartile range (IQR) (dekking2005modern) of the ensemble predictions over each trajectory, defined as
$\mathrm{IQR}=q_{0.75}-q_{0.25},$
where $q_{0.25}$ and $q_{0.75}$ denote the 25th and 75th percentiles of the ensemble predictions at a given time instant, respectively. The IQR is used instead of the standard deviation to reduce sensitivity to a small number of outlier models (dekking2005modern).
@@
In Figs. 10(a) and 10(b), we present the results for the interpolation trajectory for the heavy-phase height and the DPZ height, respectively. While the two-stage trained PINN ensemble tracks the averaged heights with good accuracy, the VNN ensemble mean overshoots with regard to the DPZ height. For both model types, the non-pretrained versions yield noticeably poorer predictions, with the effect being especially pronounced for the VNN. Moreover, we see that the mechanistic model predictions diverge, predicting flooding of the separator with the dispersed-phase at around $\tau=$380\text{\,}\mathrm{s}\text{/}$\$.
In Figs. 11(a) and 11(b), we present the extrapolation trajectory results. The overall trends are similar to those observed for the interpolation trajectories: The two-stage trained PINN ensemble provides a fairly accurate tracking (albeit a slight overshoot towards the end), the VNN and the non-pretrained models perform significantly worse and the mechanistic model predicts flooding at around $\tau=$420\text{\,}\mathrm{s}\text{/}$\$.
@@
Figure 10: Simulation results for the interpolation test trajectory. Predictions from an individual model are shown in transparent red, whereas the ensemble mean estimation is shown with a red dashed line. The reported root mean square error (RMSE) values quantify the deviation between the ensemble mean prediction of each model and the corresponding experimental trajectory over the entire trajectory. Time-averaged IQR values are reported to quantify the ensemble spread.
@@
Figure 11: Simulation results for the extrapolation test trajectory. Predictions from an individual model are shown in transparent red, whereas the ensemble mean estimation is shown with a red dashed line. The reported RMSE values quantify the deviation between the ensemble mean prediction of each model and the corresponding experimental trajectory over the entire trajectory. Time-averaged IQR values are reported to quantify the ensemble spread.